a mother is presently 5 years older than twice her daughter's sge. 10 years ago the product of the mother and daughter's age was equal to 125.What is the mom's age?
Answer by stanbon(75887)

a mother is presently 5 years older than twice her daughter's sge. 10 years ago the product of the mother and daughter's age was equal to 125.What is the mom's age?
-----------
Equations:
m = 2d + 5
(m-10)(d-10) = 125
-------------
Substitute for "m" and solve for "d:
(2d+5-10)(d-10) = 125
--------------
(2d-5)(d-10) = 125
2d^2-25d+50-125 = 0
---
2d^2 - 25d -75 = 0
-----
(d-15)(2d+5) = 0
d = 15 (daughter's age)
---
m = 2d+5 = 35 (mom's age)
===============================
Cheers,
Stan H.